If <1 and <2 are complementary angles, and m<1=5x-9 and m<2=4x, find the measure of the two angles
Answer by jrfrunner(365)

To be complementary means that the angles add up to 90 degrees
therefore <1 + <2=90 also since m<1=5x-9 and m<2=4x
then 5x-9 + 4x =90 and 9x-9=90 and x=99/9=11 so plug x=11 in 5x-9 and 4x to get your answer